OpenStudy (anonymous):

Find the area bounded by y = e^-2x, x=0, x=ln2, and the x-axis.

OpenStudy (dumbcow):

\[A =\int\limits_{0}^{\ln 2} e ^{-2x}dx\]

OpenStudy (dumbcow):

use u substitution to integrate

OpenStudy (anonymous):

I set it up just like that

OpenStudy (dumbcow):

u=-2x du = -2dx A = -1/2integral of e^u = -1/2*e^u

OpenStudy (anonymous):

ln(0) doesn't exist.